The voule of a glass vessel is 1000 cc at 20°C.What voulem of mercury should be poured into it at this temperature so that the volumeof the remaining space does not change with temperature? Coefficients of cubical expansion of mercury and glass are 1.8x10^-4/°C and 9x10^-6/°C respectively.



Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

volume of remaining space = 1000(1+T(9*10^-6)) - v(1+T(1.8*10^-4)) =V=constant w.r.t. temperature

thus, dV/dT = 0, implies 1000*9*10^-6 = v*1.8*10^-4

therefore, the answer is v = 50cc of mercury